OpenStudy (anonymous):

will give medal : ) ! plz help !! Jeremiah is making a shake by mixing two different protein powders, measured in ounces. The strawberry-flavored powder has 4 grams of protein per ounce. The banana-flavored powder has 3 grams of protein per ounce. He wants the drink to have a total of 6 ounces of powder and contain 22 grams of protein. Create a system of linear equations to represent the situation that will determine the exact number of ounces needed for each power. Based on the system you wrote, which algebraic method of solving systems of equations will you choose to solve the system? Explain your choice. Solve the system algebraically, showing your work. Then interpret the solution by explaining what your solution represents in the context of the problem.

OpenStudy (anonymous):

lets to it the easy way if you put \(x\) as the number of ounces of 4 gram protein and \(y\) as the number of ounces of \(3\) gram protein, then you know two equations \[x+y=6\\4x+3y=22\]

OpenStudy (anonymous):

solve for \(y\) in the first equation and get \(y=6-x\) then substitute it in to the second equation and get \[4x+3(6-x)=22\] and solve that one for \(x\)

OpenStudy (anonymous):

how do you get y=6 when you dived 22 by 3 which equals 7.33333333333333